500 bags of onions were purchased at $2.53 per bag. The desired markup is 49% based on selling price, but 20% spoilage is expected. What should the selling price per bag be (in $ per bag)? (Round your answer to the nearest cent.)
Answer #1

Cost of 500 bags of onion = 500*2.53=$1265

Spoilage of 20% =500*20%=100 bags, which means only 400 bags left to sell

Desired profit = 49 percent

Profit percentage = (400*S.P - 1265)/1265

0.49 = (400*S P - 1265)/1265

SP =$4.71

Hence, selling price per bag will be $4.71